OpenStudy (meilendurcer):

Which ratio is equivalent to 3/15?

OpenStudy (anonymous):

3:15 or 3 to 15

OpenStudy (anonymous):

3/15=1/5=2/10=4/20=5/25=...=(n*1)/(n*5)